How much strength of the magnetic field is necessary to produce a difference in the energy of two spin states of an electron of 10^-26J?


1
Expert's answer
2021-04-08T12:15:46-0400

∆Ep= gpβN ×2.3487 = 7×10^-26 J

∆Ee = geβe×0.34T = 6 ×10^-24J

The value of KT at 300K is 1.5033 ×10^-23 × 30= 4.51 × 10^-21

= 10^-26/4.51^-21

= 4.51 × 10^-4 T